If x + y = -23 and x - y = 7 then 2x - y equals
A.-31 B.-1 C.1 D.8 E.15

Answer by richwmiller(17219)

x+y=-23
x-y=7
2x=-16
x=-8
y=-15
-16+15=-1
B)-1